a letter is selected at random from the word HOLLYWOOD
a)What is the sample space for the experiment
b) Describe the event the letter chosen is a vowel as a subset of the sample space.
I'm really at a loss :(

Answer by AnlytcPhil(1806)

a letter is selected at random from the word HOLLYWOOD
a)What is the sample space for the experiment
HOLLYWOOD contains two L's and 3 O's but only 1 each of the letters H,Y,W,D.

So the sample space = {D,H,L,O,W,Y} 

[But this is not an 'equiprobable' sample space. 

The probability of selecting a D is 1/9, written P(D)=1/9. 
The probability of selecting a H is 1/9, written P(D)=1/9.
The probability of selecting a L is 2/9, written P(D)=2/9.
The probability of selecting a O is 3/9=1/3, written P(D)=1/3.
The probability of selecting a W is 1/9, written P(D)=1/9.
The probability of selecting a Y is 1/9, written P(D)=1/9.]

b) Describe the event the letter chosen is a vowel
as a subset of the sample space.
"Y" is a vowel in "HOLLYWOOD", so 

the event the letter chosen is a vowel = {O,Y}

Edwin